OpenStudy (anonymous):

rationalize the denominator and simplify (√:4(3s^3))/(√4:(2t^5))

OpenStudy (mandre):

Can you use the Equation tool to make it a bit more clear?

OpenStudy (anonymous):

@Mandre ok

OpenStudy (anonymous):

\[\frac{ \sqrt[4]{3s^3} }{ \sqrt[4]{2t^5} }\]

OpenStudy (anonymous):

@Mandre

OpenStudy (mandre):

The best I can do is: \[\frac{\sqrt[4]{3s ^{3}}.\sqrt[4]{2t ^{5}} }{ \sqrt[4]{2t ^{5}}.\sqrt[4]{2t ^{5}} }\] Then you get: \[\frac{\sqrt[4]{3s ^{3}.{2t ^{5}}} }{ 2t ^{5} }\] \[=\frac{ \sqrt[4]{6s ^{3}t ^{5}} }{ 2t ^{5} }\]

OpenStudy (anonymous):

thank you @Mandre

OpenStudy (luigi0210):

@Mandre kind of made a mistake

OpenStudy (luigi0210):

You have to multiply by the reciprocal not the same denominator

OpenStudy (anonymous):

@Luigi0210 ok I will post up anther question can you help me

OpenStudy (luigi0210):

I'll try